| Operation Summary for November 18, 2011 - November 21, 2011 |
|
NuMI reported receiving 9.22E18 protons on target for the period from 11/14/11 to 11/21/11. Friday November 18 The day shift began with Operations monitoring the Main Injector (MI), Linac, and Booster, with Neutron Therapy Facility (NTF) personnel preparing to treat one patient later in the morning, and with the MiniBooNE, NuMI, and the Meson FTBFT-978 experiments taking beam. During the day shift, at 8:15 AM, a Booster RF station (BRF1) tripped off; it reset. At 11:21 AM, a Linac RF station (LRF2) tripped off; it reset. At 2:06 AM, the NuMI beam permit tripped off for the first of two times in about 80 minutes due to an RF cavity (MIRF16) short reading an indeterminate state; RF personnel cycled the cavity short. Operators started a Muon ring study (formerly the Antiproton Source) at 3:33 PM. During the evening shift, Operations disable the T-978 beam at 7:08 PM. Operators ended the Muon ring study at 8:24 PM. At 9:11 PM, the NuMI beam permit tripped off due to a MIRF16 cavity short reading an indeterminate state; it reset. Saturday November 19 The midnight shift began with Operations sending beam to NuMI and MiniBooNE. During the midnight shift, at 2:03 AM, two I- Source vacuum pumps tripped off; they reset. At 4:02 AM, the NuMI beam permit tripped off due to a MIRF16 cavity short reading an indeterminate state; it reset. At 4:03 AM, MIRF15 tripped off on gap faults; it reset. At 5:54 AM, BRF17 tripped off; it reset. Operations established beam to T-978 at 7 AM. Operators started a Muon ring study around 7:06 AM. At 7:21 AM, BRF1 tripped off for the first of eight times in 53 minutes due to various problems; it reset. During the day shift, at 8:04 AM, BRF1 tripped off; it reset. At 8:34 AM, LRF1 tripped off; it reset. At 9:26 AM, BRF1 tripped off for the first of three times in ten minutes; it would not reset after the last trip. Operators bypassed the station and called in RF personnel. At 11:25 AM, the Booster East Anode Power Supply tripped off on an overload indication; it reset. During the evening shift, at 4:16 PM, BRF1 tripped off; it reset. At 5:06 PM, LRF2 tripped off and would not reset; operators called in a Linac technician. At 5:43 PM, operators reported that they had reset BRF1 four times due to 480 Volt 3- Phase absent indication; it reset. Operations disabled the T-978 beam around 7 PM. At 7:11 PM, the NuMI beam permit tripped off on a RF gap incomplete indication; operators reported that all MI RF stations were up and running and reset the trip. Operators conducted a 750 KeV line tune at 8:15 PM. Operators conducted a Booster injection tune at 8:16 PM. Operators ended the Muon ring study at 9:07 PM. MIRF15 tripped off at 9:22 PM; it reset. At 9:38 PM, BRF1 tripped off for the first of four times in about an hour and 20 minutes; it reset. Sunday November 20 The midnight shift began with Operations sending beam to NuMI and MiniBooNE. During the midnight shift, at 12:38 AM, an I- Source vacuum pump tripped off; it reset. At 12:43 AM, a NuMI BPM (NBPM6N) stopped polling; operators rebooted the BPM. At 4:40 AM, BRF1 tripped off for the first of three times in 23 minutes; it reset. Operators started a Muon ring study at 7:52 AM. Operations established beam to T-978 at 7:57 AM. During the day shift, at 10:49 AM, BRF1 tripped off; it reset. During the evening shift, Operators reported at 6:05 PM that they were unable to send changes to MECAR; operators will hold off on a reboot until the Muon ring studies end. At 6:51 PM, a Linac Klystron RF station (KRF1) tripped off on a bad PFN current indication; it reset. Operations disabled the T-978 beam at 7:24 PM. At 7:36 PM, the NuMI beam permit tripped off on a RF gap incomplete indication; all MI RF stations were up and running. Operators reset the trip. Operators ended the Muon study at 8:03 PM. Operators halted MI beam at 9:08 PM to reboot MECAR. At 9:24 PM, BRF1 tripped off for the fist of four times in ten minutes; it reset. Monday November 21 The midnight shift began with Operations sending beam to NuMI and MiniBooNE. During the midnight shift, at 2:46 AM, BRF1 tripped off; it reset. At 3:14 AM, operators reported many consecutive trips of BRF1; operators adjusted the anode voltage. At 3:15 AM, NBPM65 stopped polling; operators rebooted it. Operations established beam to T-978 at 7:05 AM. At 7:15 AM, two I- Source vacuum pumps tripped off; they reset. Operations Coordinator (OpsCo) The OpsCo approved of the plans by Booster and MI personnel to conduct studies. Linac personnel will take 90-minutes of downtime this morning to work on KRF1. During this time NuMI personnel will access their Absorber room to inspect a water skid. |
|
More Information Comments and Suggestions |
| last modified 11/21/2011 email Fermilab |
| Security, Privacy, Legal | Use of Cookies | |